# Partial fraction decompositions on hyperplane arrangements

### v1

@@ 3. The braid arrangement
We demonstrate our primary decomposition (Corollary 2.9) on braid arrangements, which form an important family of hyperplane arrangements arising across algebra, geometry, topology, and combinatorics (see, e.g., [Fei]). Braid arrangements also appear in physics via denominators of broken Parke–Taylor functions [Fro] and Parke–Taylor forms [BEPV].
@@
Table 1. The first column lists all the partitions of 55, the second states the number of flats associated with a given partition, the third gives the number of hyperplanes for each such flat, and the fourth gives an example of such a flat. We use ijij to denote the polynomial $x_{i}-x_{j}$ for brevity.
@@
Let $r=5$ and $d=8$, and consider the braid arrangement $\mathcal{B}_{5}$. There are $n=\binom{5}{2}=10$ linear forms in our arrangement $L$. Then Corollary 2.9 gives a primary decomposition of $I_{L,d}$ as an intersection of the following kinds of ideals:
@@
  • five ideals, each to the power of 44, corresponding to flats associated with $(4,1)$;
  • ten ideals, each to the power of 22, corresponding to flats associated with $(3,2)$;
  • ten ideals, each to the power of 11, corresponding to flats associated with $(3,1,1)$.
One may notice from Table 1 and Example 3.2 that if $\lambda\succeq\mu$ under the dominance order on partitions (recall that $\lambda\succeq\mu$ under the dominance order if and only if $\lambda_{1}+\cdots+\lambda_{i}\geq\mu_{1}+\cdots+\mu_{i}$ for all $i\in[r]$), then the size of a flat associated with $\lambda$ is larger than that of $\mu$. The following result shows that this holds more generally, thus showing that the set of partitions associated with flats appearing in our primary decomposition (Corollary 2.9) is a filter in the poset of partitions.
Proposition $3.$
Suppose $\lambda,\mu\vdash r$ be partitions, with $\lambda\succeq\mu$ under the dominance order. Then $|S_{\lambda}|\geq|S_{\mu}|$, where $S_{\lambda}$ (resp. $S_{\mu}$) denotes any flat associated with $\lambda$ (resp. $\mu$). In particular, if flats corresponding to the partition $\mu$ appear in the primary decomposition of Corollary 2.9, then so do flats corresponding to the partition $\lambda$.
@@
Let $\lambda=(\lambda_{1},\ldots,\lambda_{r})$ and $\mu=(\mu_{1},\ldots,\mu_{r})$, with $\lambda_{i},\mu_{i}\geq 0$ for all ii. It suffices to assume that $\lambda\succeq\mu$ is a cover relation. By [Bry, Proposition 2.3], this implies that there exists $i,j\in[r]$ with $i<j$ such that $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$, $\lambda_{i}-\lambda_{j}\geq 2$, and $\lambda_{k}=\mu_{k}$ for all $k\neq i,j$. Then $|S_{\lambda}|\geq|S_{\mu}|$ if and only if:
@@ 4. Adjoint polynomials
A rapidly developing field at the intersection of physics, combinatorics, and algebraic geometry is the study of positive geometries and canonical forms [Lam, RST]. Central mathematical objects in this study are projective polytopes and their adjoint polynomials. In this section, we highlight how the primary decomposition in Corollary 2.9 can be used to easily deduce vanishing properties of adjoint polynomials. We define this polynomial, and refer the reader to the paper [Gae] for further introduction to the relevant terminology.
@@
The form of $\mathrm{adj}_{C}$ is independent of the choice of triangulation of $C$, and $\mathrm{adj}_{C}$ is homogeneous and of degree $|V(C)|-\dim(C)$ [Gae, War]. Furthermore, Kohn–Ranestad showed that the residual arrangement of the hyperplane arrangement defining the facets of the dual cone $C^{\vee}$ must vanish on the adjoint polynomial; the adjoint polynomial is also uniquely characterized by certain vanishing conditions on this residual arrangement [BW, KR].
The above definition of adjoint polynomial makes it clear that adjoint polynomials are in ideals of the form $I_{L,d}$, where in this case the hyperplanes defining the ideal are defined by the $\ell_{v}$ from Definition 4.1. Consequently, using the primary decomposition of Corollary 2.9, one can determine certain properties about where adjoint polynomials vanish, without needing to compute any of the $a_{S}$ coefficients or the residual arrangement of the dual cone. Corollary 2.9 implies the following vanishing behavior of adjoint polynomials.
@@
which, using Corollary 2.9, can be expressed as the primary decomposition:
@@ 5. Partial fraction decompositions and algorithmic aspects
Generally, the least common multiple of all of the terms in the PFD expression is exactly $\ell_{1}\cdots\ell_{n}$. By putting all of the terms of the PFD under this common denominator, we see that such a PFD exists exactly when the polynomial $f$ can be expressed as an $R$-linear combination of $d$-fold products of the $\ell_{i}$’s. In other words, a PFD of degree $d$ exists for $f/(\ell_{1}\cdots\ell_{n})$ if and only if $f$ lies in the ideal $I_{L,d}$ studied in §2. We can thus interpret our results on primary decompositions of $I_{L,d}$ as vanishing criteria for existence of PFDs. We first consider the case of hyperplane arrangements in linearly general position.
@@
We now turn to the case of arbitrary hyperplane arrangements, both in the projective and affine settings. The following results on PFDs follow from Corollary 2.9 and Theorem 2.12.
@@ 6. Towards physics
In this section, we apply our algorithm from §5 to compute PFDs of rational functions arising from two different contexts in physics: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). The code for these examples can be found here.
@@ 6.1. Feynman integrals
In preparation for the approaching upgrade to the Large Hadron Collider [AAA+], much work has been done in developing efficient algorithms to aid the computation of Feynman integrals [HvM, BFP, CGM]. Many of these algorithms are based on integration-by-parts (IBP) reductions, in which a Feymann integral is decomposed as a linear combination of master integrals (MIs). These MIs are a choice of basis for the family of Feynman integrals under consideration. The coefficients of the linear combination are rational functions in the external kinematics and the dimensional regularization parameter $\epsilon$. These rational functions are often very large, and can make calculations cumbersome. The authors of [BW+] suggest storing the coefficients of IBP reductions in the form of a PFD, and introduce their own PFD algorithm as a “powerful method to reduce the byte size of the analytic IBP reduction coefficients”. Their paper includes an application to IBP reduction coefficients arising from the Feynman integral associated with the the two-loop five-point non-planar double pentagon diagram shown in Figure 2. These coefficients were calculated (with respect to a specific choice of MIs) in [BBD+]. The authors of [HvM] suggest using these coefficients as a benchmark for testing PFD algorithms for Feynman integrals. In this section we naively apply our PFD algorithm to two of these coefficients, a “small” and a “large” example, and discuss the results. We hope this may serve as a starting point for further refinement by physicists.
@@ 6.1.1. Small IBP coefficient
This IBP coefficient is a rational function in six variables. These variables are the dimensional regulatization parameter $\epsilon$ and the symbolic Mandelstam variables $s_{12}$, $s_{15}$, $s_{23}$, $s_{34}$, and $s_{45}$ (Lorentz invariant functions of the external momenta $p_{i}$). Explicitly, the polynomial ring here is $\mathbb{Q}[s_{12},s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is a degree $11$ polynomial, and the denominator can be factorized as a product of the hyperplanes:
@@
We find that the input rational function contains no spurious denominator factors and that the largest $d$ for which the numerator is in the ideal $I_{L,d}$ is 88. This implies the existence of a PFD of degree 88 as in (7), with degree 44 denominators. Algorithm 2 produces such a PFD with $173$ terms (where the number of distinct terms is the the number of distinct denominators appearing in the PFD). These results were computed within seconds.
@@ 6.1.2. Large IBP coefficient
This IBP coefficient is a rational function in the same six variables as above. To reduce computation time, one of the Mandelstam variables can be set to 11. This variable can later be re-introduced through dimensional arguments. We choose to set $s_{12}=1$, and so the polynomial ring here is $\mathbb{Q}[s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is then a degree $30$ polynomial in five variables, and the denominator can be written as a product of the following $29$ hyperplanes:
@@
Given that we have set $s_{12}=1$, we have suppressed the $30$th denominator hyperplane $\ell_{30}=s_{12}$. Without making any modifications to Algorithm 2, the time it takes to construct the ideals $I_{L,d}$ grows exponentially with $d$, already taking hours at $d=7$. By choosing subsets of the generators of the ideals one can dramatically reduce computation time. The only drawback is that the numerator not being in a subset of $I_{L,d}$ does not rule out the non-existence of a PFD of degree $d$. Using this strategy, we have found that the numerator is in an ideal contained in the ideal $I_{L,17}$. This smaller ideal was generated using $5985$ of the $51895935$ generators of $I_{L,17}$ (we did not include $\ell_{22},\ldots,\ell_{29}$) and took $194$ seconds to construct. Testing that the numerator is in this ideal took $930$ seconds, and generating the corresponding PFD took $2217$ seconds. This PFD has $546$ terms (counted by distinct denominators) and degree $12$ denominators. Comparing with the times given in [HvM, Table 2], our algorithm implemented in Macaulay2 is competitive with other modern PFD algorithms for Feynman integrals.
@@ 6.2. Wavefunction coefficients
The function $\psi/(8Y_{14}Y_{24}Y_{34})$ is the canonical form of the cosmological polytope whose facet hyperplanes are the denominator factors. In [FPSW], the authors conjecture a formula for the PFD of flat-space wavefunction coefficients associated with arbitrary tree-level cosmological Feynman diagrams. An amended version of this conjecture was recently proven in [Dun]. That a PFD of this form exists in this case is clear from the fact that the numerator of $\psi$ (once all the terms have been collected under a common denominator) is in the ideal:
@@
(as can be checked using Macaulay2), and that the numerator is homogeneous of degree 77. And so, using the command // one can express $\psi$ in terms of the generators of this ideal, yielding:
@@
Note that this is not the PFD one arrives at for the diagram 2 using the process outlined in [Dun], though both PFDs have the same number of terms. This means that there can be multiple different PFDs of tree-level flat space wavefunction coefficients which have the same number of terms, and in which all numerators are $\pm 1$ and all denominators have the same degree. In other words, such “minimal” flat-space wavefunction coefficient PFDs are not unique.
One may also wonder about the vanishing properties of the numerator. Using the Matroids package for Macaulay2, one finds that the matroid defined by lines $\ell_{i}$ has $88$ flats of five elements, $44$ of six elements, $12$ of seven elements, 11 of eight elements, 33 of nine elements, and 33 of ten elements. Corollary 2.9 then tells us that the numerator vanishes on each of these flats $S$, with order |S|4|S|-4. For example, the flats of ten elements are:
@@ 7. Future directions and outlook
We then have the following two PFDs of degree 33:
@@
Through our study of ideals of the form $I_{L,d}$, we have proven a number of results about PFDs of rational functions on hyperplane arrangements. This led to new ways of understanding the geometry of adjoint polynomials and canonical forms, as well as a novel algorithmic approach for constructing PFDs of a tailored form. We have demonstrated how our findings have direct applications to a variety of problems in physics and hope that this paper inspires further study of PFDs by algebraists.

### v2

@@ 3. The braid arrangement
We apply the primary decomposition in Theorem 2.5 to braid arrangements, which form an important family of hyperplane arrangements arising across algebra, geometry, topology, and combinatorics (see, e.g., [Fei]). Braid arrangements also appear in physics via denominators of broken Parke–Taylor functions [Fro] and Parke–Taylor forms [BEPV].
@@
Table 1. The first column lists all the partitions of $5$, the second states the number of flats associated with a given partition, the third gives the number of hyperplanes for each such flat, and the fourth gives an example of such a flat. We use ijij to denote the polynomial $x_{i}-x_{j}$ for brevity.
@@
Let $r=5$ and $d=8$, and consider the braid arrangement $\mathcal{B}_{5}$. There are $n=\binom{5}{2}=10$ linear forms in our arrangement $L$. Then Theorem 2.5 gives a primary decomposition of $I_{L,d}$ as an intersection of the following kinds of ideals:
@@
  • five ideals, each to the power of $4$, corresponding to flats associated with $(4,1)$;
  • ten ideals, each to the power of $2$, corresponding to flats associated with $(3,2)$;
  • ten ideals, each to the power of $1$, corresponding to flats associated with $(3,1,1)$.
One may notice from Table 1 and Example 3.2 that if $\lambda\succeq\mu$ under the dominance order on partitions (recall that $\lambda\succeq\mu$ under the dominance order if and only if $\lambda_{1}+\cdots+\lambda_{i}\geq\mu_{1}+\cdots+\mu_{i}$ for all $i\in[r]$), then the size of a flat associated with $\lambda$ is larger than that of $\mu$. The following result shows that this holds more generally, thus showing that the set of partitions associated with flats appearing in the primary decomposition of Theorem 2.5 is a filter in the poset of partitions.
Proposition 3.3.
Suppose $\lambda,\mu\vdash r$ be partitions, with $\lambda\succeq\mu$ under the dominance order. Then $|S_{\lambda}|\geq|S_{\mu}|$, where $S_{\lambda}$ (resp. $S_{\mu}$) denotes any flat associated with $\lambda$ (resp. $\mu$). In particular, if flats corresponding to the partition $\mu$ appear in the primary decomposition of Theorem 2.5, then so do flats corresponding to the partition $\lambda$.
@@
Let $\lambda=(\lambda_{1},\ldots,\lambda_{r})$ and $\mu=(\mu_{1},\ldots,\mu_{r})$, with $\lambda_{i},\mu_{i}\geq 0$ for all $i$. It suffices to assume that $\lambda\succeq\mu$ is a cover relation. By [Bry, Proposition 2.3], this implies that there exists $i,j\in[r]$ with $i<j$ such that $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$, $\lambda_{i}-\lambda_{j}\geq 2$, and $\lambda_{k}=\mu_{k}$ for all $k\neq i,j$. Then $|S_{\lambda}|\geq|S_{\mu}|$ if and only if:
@@ 4. Adjoint polynomials
A rapidly developing field at the intersection of physics, combinatorics, and algebraic geometry is the study of positive geometries and canonical forms [Lam, RST]. Central mathematical objects in this study are projective polytopes and their adjoint polynomials. In this section, we highlight how the primary decomposition in Theorems 2.5 and 2.7 can be used to easily deduce vanishing properties of adjoint polynomials. We define this polynomial, and refer the reader to the paper [Gae] for further introduction to the relevant terminology.
@@
The form of $\mathrm{adj}_{C}$ is independent of the choice of triangulation of $C$, and $\mathrm{adj}_{C}$ is homogeneous and of degree $|V(C)|-\dim(C)$ [Gae, War]. Furthermore, Kohn–Ranestad showed that the adjoint polynomial must vanish on the residual arrangement of the hyperplane arrangement defining the facets of the dual cone $C^{\vee}$; the adjoint polynomial is also uniquely characterized by certain vanishing conditions on this residual arrangement [BW, KR].
The above definition of adjoint polynomial makes it clear that adjoint polynomials are in ideals of the form $I_{L,d}$, where in this case the hyperplanes defining the ideal are defined by the $\ell_{v}$ from Definition 4.1. Consequently, using the primary decomposition of Theorem 2.5, one can determine certain properties about where adjoint polynomials vanish, without needing to compute any of the $a_{S}$ coefficients or the residual arrangement of the dual cone. Theorem 2.5 implies the following vanishing behavior of adjoint polynomials.
@@
which, using Theorem 2.5, can be expressed as the primary decomposition:
@@ 5. Partial fraction decompositions and algorithmic aspects
Typically, the least common multiple of all of the terms in the PFD expression is exactly $\ell_{1}\cdots\ell_{n}$. By putting all of the terms of the PFD under this common denominator, we see that such a PFD exists exactly when the polynomial $f$ can be expressed as an $R$-linear combination of $d$-fold products of the $\ell_{i}$’s. In other words, a PFD of degree $d$ exists for $f/(\ell_{1}\cdots\ell_{n})$ if and only if $f$ lies in the ideal $I_{L,d}$ presented in §2. We can thus interpret the primary decompositions of $I_{L,d}$ as vanishing criteria for existence of PFDs. We first consider the case of hyperplane arrangements in linearly general position.
@@
We now turn to the case of arbitrary hyperplane arrangements, both in the projective and affine settings. The following results on PFDs follow from Theorems 2.5 and 2.7.
@@ 6. Towards physics
In this section, we apply our algorithm from §5 to compute PFDs of rational functions arising from two different contexts: Feynman integrals (§6.1) and flat-space wavefunction coefficients (§6.2). The code for these examples can be found here.
@@ 6.1. Feynman integrals
In preparation for the approaching upgrade to the Large Hadron Collider [AAA+], much work has been done in developing efficient algorithms to aid the computation of Feynman integrals [HvM, BFP, CGM]. Many of these algorithms are based on integration-by-parts (IBP) reductions, in which a Feymann integral is decomposed as a linear combination of master integrals (MIs). These MIs are a choice of basis for the family of Feynman integrals under consideration. The coefficients of the linear combination are rational functions in the external kinematics and the dimensional regularization parameter $\epsilon$. These rational functions are often very large, and can make calculations cumbersome. The authors of [BW+] suggest storing the coefficients of IBP reductions in the form of a PFD, and introduce their own PFD algorithm as a “powerful method to reduce the byte size of the analytic IBP reduction coefficients”. Their paper includes an application to IBP reduction coefficients arising from the Feynman integral associated with the two-loop five-point non-planar double pentagon diagram shown in Figure 2. These coefficients were calculated (with respect to a specific choice of MIs) in [BBD+]. The authors of [HvM] suggest using these coefficients as a benchmark for testing PFD algorithms for Feynman integrals. In this section we naively apply our PFD algorithm to two of these coefficients, a “small” and a “large” example, and discuss the results. We hope this may serve as a starting point for further refinement.
@@ 6.1.1. Small IBP coefficient
This IBP coefficient is a rational function in six variables. These variables are the dimensional regularization parameter $\epsilon$ and the Mandelstam variables $s_{12}$, $s_{15}$, $s_{23}$, $s_{34}$, and $s_{45}$ (Lorentz invariant functions of the external momenta $p_{i}$ shown in Figure 2). Explicitly, the polynomial ring here is $\mathbb{Q}[s_{12},s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is a degree $11$ polynomial with $785$ terms. This polynomial can be accessed through the links provided in [BBD+]. The denominator can be factorized as a product of the hyperplanes:
@@
We find that the input rational function contains no spurious denominator factors and that the largest $d$ for which the numerator is in the ideal $I_{L,d}$ is $8$. This implies the existence of a PFD of degree $8$ as in (7), with degree $4$ denominators. Algorithm 2 produces such a PFD with $173$ terms (where the number of distinct terms is the number of distinct denominators appearing in the PFD). These results were computed within seconds.
@@ 6.1.2. Large IBP coefficient
This IBP coefficient is a rational function in the same six variables as above. To reduce computation time, one of the Mandelstam variables can be set to $1$. This variable can later be re-introduced through dimensional arguments. We choose to set $s_{12}=1$, and so the polynomial ring here is $\mathbb{Q}[s_{15},s_{23},s_{34},s_{45},\epsilon]$. The numerator is then a degree $30$ polynomial in five variables with $73763$ terms. As before, it can be accessed through the links in [BBD+]. The denominator can be written as a product of the following $29$ hyperplanes:
@@
Given that we have set $s_{12}=1$, we have suppressed the $30$th denominator hyperplane $\ell_{30}=s_{12}$. Without making any modifications to Algorithm 2, the time it takes to construct the ideals $I_{L,d}$ grows exponentially with $d$, already taking hours at $d=7$. By choosing subsets of the generators of the ideals one can dramatically reduce computation time. The only drawback is that the numerator not being in a subset of $I_{L,d}$ does not rule out the non-existence of a PFD of degree $d$. Using this strategy, we have found that the numerator is in an ideal contained in the ideal $I_{L,17}$. This smaller ideal was generated using $5985$ of the $51895935$ generators of $I_{L,17}$ (we did not include $\ell_{22},\ldots,\ell_{29}$) and took $194$ seconds to construct. Testing that the numerator is in this ideal took $930$ seconds, and generating the corresponding PFD took $2217$ seconds. This PFD has $546$ terms (counted by distinct denominators) and degree $12$ denominators. Comparing with the times given in [HvM, Table 2], our algorithm implemented in Macaulay2 is competitive with other state-of-the-art PFD algorithms for Feynman integrals.
@@ 6.2. Wavefunction coefficients
The function $\psi/(8Y_{14}Y_{24}Y_{34})$ is the canonical form of the cosmological polytope whose facet hyperplanes are the denominator factors. In [FPSW], the authors conjecture a formula for the PFD of flat-space wavefunction coefficients associated with arbitrary tree-level cosmological Feynman diagrams. An amended version of this conjecture was recently proven in [Dun]. That a PFD of this form exists in this case is clear from the fact that the numerator of $\psi$, a polynomial of degree $7$ with $204$ terms (once all the terms have been collected under a common denominator), is in the ideal:
@@
(as can be checked using Macaulay2), and that the numerator is homogeneous of degree $7$. And so, using the command // one can express $\psi$ in terms of the generators of this ideal, yielding:
@@
Note that this is not the PFD one arrives at for Figure 2 using the process outlined in [Dun], though both PFDs have the same number of terms. This means that there can be multiple different PFDs of tree-level flat space wavefunction coefficients which have the same number of terms, and in which all numerators are $\pm 1$ and all denominators have the same degree. In other words, such “minimal” flat-space wavefunction coefficient PFDs are not unique; this may be related to the different triangulations of cosmological polytopes found in [BDFJ].
One may also wonder about the vanishing properties of the numerator. Using the Matroids package for Macaulay2, one finds that the matroid defined by lines $\ell_{i}$ has $88$ flats of five elements, $44$ of six elements, $12$ of seven elements, $1$ of eight elements, $3$ of nine elements, and $3$ of ten elements. Theorem 2.5 then tells us that the numerator vanishes on each of these flats $S$, with order |S|4|S|-4. For example, the flats of ten elements are:
@@ 7. Future directions and outlook
We then have the following two PFDs of degree $3$:
@@
By recasting the existence of PFDs as an ideal membership problem, we have proven a number of results about PFDs of rational functions on hyperplane arrangements. This led to new ways of understanding the geometry of adjoint polynomials and canonical forms, as well as a novel algorithmic approach for constructing PFDs of a tailored form. We have demonstrated how our findings have direct applications to a variety of problems in high-energy physics and hope that this paper inspires further study of PFDs by algebraists.